Express the following equations of a parabola in standard form and in each case state the coordinates of its vertex,focus and the ends of the latus rectum.

(a) x^2− 2x − 4y = 0,

(b) y^2 + 12x − 48 = 0,

(c) x^2 − 6x − 2y + 7 = 0.


1
Expert's answer
2021-07-22T11:13:54-0400

(a)


"x^2-2x+1=4y+1"

"(x-1)^2=4\cdot1(y+\dfrac{1}{4})"

"p=1>0"


"h=1, k=-\dfrac{1}{4}"


"k+p=-\dfrac{1}{4}+1=\dfrac{3}{4}"

Vertex: "V(1, -\dfrac{1}{4})"


Focus: "F(1, \dfrac{3}{4})"

Latus rectum: "y=\dfrac{3}{4}"


"x^2-2x+1=4(\dfrac{3}{4})+1"

"x^2-2x-3=0"

"(x+1)(x-3)=0"

"x_1=-1, x_2=3"

The ends of the latus rectum


"(-1, \dfrac{3}{4}), (3, \dfrac{3}{4})"


(b)


"y^2 + 12x \u2212 48 = 0"

"y^2=-4\cdot3(x-4)"

"p=-3<0"


"h=4, k=0"


"h+p=4-3=1"

Vertex: "V(4, 0)"


Focus: "F(1, 0)"

Latus rectum: "x=1"


"y^2 + 12(1) \u2212 48 = 0"

"y^2=36"

"y_1=-6, y_2=6"

The ends of the latus rectum


"(1, -6), (1, 6)"



(c)


"x^2 \u2212 6x \u2212 2y + 7 = 0"

"x^2-6x+9=2y+2"

"(x-3)^2=4\cdot(\dfrac{1}{2})(y+1)"

"p=\dfrac{1}{2}>0"


"h=3, k=-1"

"k+p=-1+\dfrac{1}{2}=-\dfrac{1}{2}"


Vertex: "V(3, -1)"


Focus: "F(3, -\dfrac{1}{2})"

Latus rectum: "y=-\dfrac{1}{2}"

"x^2 − 6x − 2(-\dfrac{1}{2}) + 7 =0"

"x^2-6x+8=0"

"(x-2)(x-4)=0"

"x_1=2, x_2=4"

The ends of the latus rectum


"(2, -\dfrac{1}{2}), (3, -\dfrac{1}{2})"